# Where Did It All Go Wrong? A Hierarchical Look into Multi-Agent Error Attribution

### v1

@@ Abstract
Error attribution in Large Language Model (LLM) multi-agent systems presents a significant challenge in debugging and improving collaborative AI systems. Current approaches to pinpointing agent and step level failures in interaction traces - whether using all-at-once evaluation, step-by-step analysis, or binary search - fall short when analyzing complex patterns, struggling with both accuracy and consistency. We present ECHO (Error attribution through Contextual Hierarchy and Objective consensus analysis), a novel algorithm that combines hierarchical context representation, objective analysis-based evaluation, and consensus voting to improve error attribution accuracy. Our approach leverages a positional-based leveling of contextual understanding while maintaining objective evaluation criteria, ultimately reaching conclusions through a consensus mechanism. Experimental results demonstrate that ECHO outperforms existing methods across various multi-agent interaction scenarios, showing particular strength in cases involving subtle reasoning errors and complex interdependencies. Our findings suggest that leveraging these concepts of structured, hierarchical context representation combined with consensus-based objective decision-making, provides a more robust framework for error attribution in multi-agent systems.
@@ 1 Introduction
Multi-Agent Systems (MASs) have demonstrated remarkable performance across coding [6], medical QA [7], and financial decision-making [8]. However, their multi-step nature makes them vulnerable to compounding errors, where early mistakes amplify through subsequent steps and can derail the entire system. As a result, identifying the initial error’s source—both agent and step—becomes crucial for mitigating such failures and improving these systems.
@@ 2.1 Evaluation of LLM Agents
Much research has focused on automated LLM evaluation across various tasks [13, 14, 15], with recent efforts shifting toward assessing the agentic capabilities of LLMs. AgentBench [16] tests task completion across diverse environments like online shopping and householding, while MLAgentBench [17] evaluates LLM agents’ abilities to perform machine learning experimentation. AssistantBench [18] features realistic, time-intensive web tasks requiring multi-turn interactions. Agent-as-a-Judge [19] leverages LLM agents to evaluate their other agent peers through tool interactions. ToolFormer [20], AnyTool [21], and ToolBench [22] benchmark LLMs’ API-calling capabilities via tool-use. As LLM agent interactions evolve, evaluation frameworks now address multi-agent scenarios: MultiAgentBench [23] measures coordination and competitive dynamics, while SwarmBench [24] assesses swarm intelligence through coordination tasks like foraging and flocking.
@@ 2.2 Error Attribution
While LLM evaluation typically measures task completion and output quality, error attribution focuses on systematically classifying error types and patterns to understand fundamental failure modes [25, 26, 27, 28]. ReaLMistake [29] assesses LLMs’ error attribution capabilities in generated output, while SynCheck [30] leverages decoding dynamics to verify sentence trustworthiness and backtracks the unfaithfulness of generated sentences. Self-Backtracking [31] trains LLMs to localize and correct their own reasoning errors through supervised fine-tuning. Similarly, Process Reward Models (PRMs) [32] evaluate intermediate reasoning step correctness through process annotation, while ProcessBench [33] measures PRMs’ step-wise evaluation capabilities. These approaches focus solely on single-agent error analysis, whereas the Who&When dataset [9] extends error attribution to multi-agent settings by identifying both the responsible agent and erroneous step.
@@ 3.1 Hierarchical Context Representation
Hierarchical representation operates at 4 levels $L_{1}$ through $L_{4}$, as seen in Appendix A.1 and A.3, to extract key information from agent / step interactions (via regex pattern matching). While LLMs could potentially be used for more flexible content extraction, the computational overhead would be significant - a trade-off we explore further in Section 4.3. Its implementation employs specialized content extraction mechanisms for each layer $C_{i}$, as seen below:
@@ 3.2 Objective Analysis
This approach prevents echo-chamber effects in attribution through intentional perspective diversity. While this was implemented in ECHO through role-specific prompts and temperatures, the framework accommodates other diversification methods (different models, LLM configurations, or reasoning strategies) - all serving the core principle of bias mitigation through analytical diversity.
@@ 3.3 Consensus Voting
The consensus mechanism processes 3 key components from each analysis $A_{t}^{j}$ (where $t\in\{\text{agent},\text{step}\}$): (1) Primary Conclusions - that considers core attribution decisions including the attribution type, responsible agent(s), and specific error step; (2) Agent Evaluations - that provides assessments of each agent’s error likelihood and supporting evidence; and (3) Alternative Hypotheses - that considers secondary attribution possibilities that maintain reasonable likelihood.
@@ 4.2.1 Performance of ECHO
Step-Level Attribution: Although ECHO still outperforms prior methods with statistical significance, step-level precision still proves challenging, with exact matches at 27–28%. However, accuracy improves significantly when accounting for step tolerance. With ±3 steps, accuracy reaches 42.1% for hand-crafted dataset with ground truth; and when further extended to ±5 steps, accuracy reaches 61.4% - again, maintaining superior performance over baselines.
@@ 4.3 Impact of ECHO Components Via Ablation
Hierarchical context shows clear advantages over fixed context windows within the context-aware framework. Shifting from fixed ±1 step (I1) to hierarchical context (I2) yields significant improvements in hand-crafted datasets: +16.1% in agent-level accuracy (P < 0.05) and +1.9% in step-level accuracy (not significant), as shown in Table 2. A similar trend is exhibited for the algorithm-generated dataset, with a +6.2% in agent-level accuracy and +3.8% in step-level accuracy. This validates graduated context preservation, where detail decreases with distance from target steps. While analysis for the hand-crafted dataset was limited to shorter traces due to computational constraints with context-aware agents, the consistent accuracy improvements highlight hierarchical context’s value - a benefit fully realized when combined with objective analysis.
@@ A.2 Fixed-Window Context
␣␣␣␣Extractagentcontextsfromconversationhistory.
␣␣␣␣Eachcontextincludesthepreviousagent,currentagent,andnextagent.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Listoftuplescontaining(prev_agent,current_agent,next_agent)foreachagent
@@ A.3 Hierarchical Context Extraction
␣␣␣␣Extractkeydecisionormainpointfromagentcontentusingregexpatterns.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedkeydecision
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Keydecisionormainpointfromtheagent’scontent
@@
␣␣␣␣Createabriefsummaryofagentcontentusingregexpatterns.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsinthesummary
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Briefsummaryoftheagent’scontent
@@
␣␣␣␣Extractmilestone-basedinformationfromagentcontentusingregexpatterns.
␣␣␣␣Thisprovidesahigherlevelofabstractionthanbriefsummariesfordistantcontexts.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣agent_content:Thefullcontentoftheagent
␣␣␣␣␣␣␣␣max_words:Maximumwordsintheextractedmilestones
␣␣␣␣␣␣␣␣context_type:Typeofcontexttofocuson(handoff,decision_quality,error_propagation,general)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Milestone-basedinformationfromtheagent’scontent
@@
␣␣␣␣Extracthierarchicalagentcontextsfromconversationhistory.
␣␣␣␣Usesgraduateddetaillevelsbasedondistancefromcurrentagent.
␣␣␣␣Args:
␣␣␣␣␣␣␣␣conversation_history:Listofconversationturnswithagentinformation
␣␣␣␣␣␣␣␣dataset_name:Nameofthedatasetbeingprocessed(affectshowagentinfoisextracted)
␣␣␣␣Returns:
␣␣␣␣␣␣␣␣Listofdictionariescontaininghierarchicalcontextforeachagent
@@ A.4 Context Step Aware Agent
␣␣␣␣Context-AwareStepAgentthatanalyzesanagentinthecontextofitspreviousandnextagents
␣␣␣␣toarguewhytheerrorhappenedinthisagent’sstep.
@@
␣␣␣␣␣␣␣␣InitializetheContext-AwareStepAgent.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣model_id:ThemodelIDtousefortheagent
␣␣␣␣␣␣␣␣␣␣␣␣temperature:Thetemperaturetouseforgeneration
@@
␣␣␣␣␣␣␣␣YouareaContext-AwareStepAgentthatanalyzesanagent’sactionsinthecontextofthepreviousandnextagents.
␣␣␣␣␣␣␣␣YourtaskistoarguewhytheerrorhappenedinYOURagent’sstep.
␣␣␣␣␣␣␣␣Yourtask:
␣␣␣␣␣␣␣␣1.Analyzewhatinformationwasreceivedbyyouragentfromthepreviousagent(ifany)
␣␣␣␣␣␣␣␣2.Analyzewhatinformationwasgeneratedbyyouragent
␣␣␣␣␣␣␣␣3.Analyzehowyouragent’soutputaffectedthenextagent(ifany)
␣␣␣␣␣␣␣␣4.MakeastrongargumentforwhyYOURAGENTcausedthefinalwronganswer,usingthegroundtruthasevidence
␣␣␣␣␣␣␣␣Input:
␣␣␣␣␣␣␣␣-GroundTruth:[GROUND_TRUTH]
␣␣␣␣␣␣␣␣-FinalAnswer:[FINAL_ANSWER]
␣␣␣␣␣␣␣␣-AgentContext:Informationabouttheprevious,current,andnextagents
␣␣␣␣␣␣␣␣Outputyourresponsewiththefollowingclearsectionheaders:
␣␣␣␣␣␣␣␣##Purpose:
␣␣␣␣␣␣␣␣Describethepurposeofthisagentstep-whatwasthisagenttryingtoaccomplish?
␣␣␣␣␣␣␣␣##AssumptionsandInformation:
␣␣␣␣␣␣␣␣Listtheassumptionsandinformationthisagentwasgivenfromthepreviousagentorcontext.
␣␣␣␣␣␣␣␣##Errors:
␣␣␣␣␣␣␣␣Describewhatthisagentdidwrong(ifanything).Bespecificaboutanymistakes,misunderstandings,orincorrectreasoning.
␣␣␣␣␣␣␣␣##Evidence:
␣␣␣␣␣␣␣␣Provideevidencefromthegroundtruththatsupportsyourerrorattribution.Explainhowthisagent’sactionsdirectlyledtothewrongfinalanswer.
␣␣␣␣␣␣␣␣Remember:YoumustarguethatYOURagentcausedtheerror.Bepersuasiveanduseevidence.
@@
␣␣␣␣␣␣␣␣Analyzeanagentinthecontextofitspreviousandnextagentsandgenerateanargument
␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
␣␣␣␣␣␣␣␣␣␣␣␣prev_agent:Thepreviousagent(orNoneiffirstagent)
␣␣␣␣␣␣␣␣␣␣␣␣current_agent:Thecurrentagentbeinganalyzed
␣␣␣␣␣␣␣␣␣␣␣␣next_agent:Thenextagent(orNoneiflastagent)
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
@@
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AgentContext:
␣␣␣␣␣␣␣␣{agent_context}
␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthecontextofthepreviousandnextagents,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
@@
␣␣␣␣␣␣␣␣Analyzeanagentusinghierarchicalcontextandgenerateanargument
␣␣␣␣␣␣␣␣forwhythisagentcausedtheerror.
␣␣␣␣␣␣␣␣Args:
␣␣␣␣␣␣␣␣␣␣␣␣step_id:TheIDofthestep(e.g.,"step_1")
␣␣␣␣␣␣␣␣␣␣␣␣hierarchical_context:Dictionarycontaininghierarchicalcontextinformation
␣␣␣␣␣␣␣␣␣␣␣␣ground_truth:Thegroundtruthanswer
␣␣␣␣␣␣␣␣␣␣␣␣final_answer:Thefinalanswergiven
␣␣␣␣␣␣␣␣␣␣␣␣query:Theoriginalquery/question
␣␣␣␣␣␣␣␣Returns:
␣␣␣␣␣␣␣␣␣␣␣␣JSONargumentforwhythisagentcausedtheerror
@@
␣␣␣␣␣␣␣␣OriginalQuery:{query}
␣␣␣␣␣␣␣␣{ground_truth_section}
␣␣␣␣␣␣␣␣FinalAnswer:{final_answer}
␣␣␣␣␣␣␣␣AgentContext:
␣␣␣␣␣␣␣␣{agent_context}
␣␣␣␣␣␣␣␣Pleaseanalyzethisagentinthehierarchicalcontextoftheentireconversation,andprovideastrongargumentforwhyTHISagentcausedthefinalwronganswer.
␣␣␣␣␣␣␣␣Usethesectionheadersspecifiedinyourinstructions(Purpose,AssumptionsandInformation,Errors,Evidence).
␣␣␣␣␣␣␣␣Note:Younowhaveaccesstothefullconversationcontextatdifferentdetaillevels:
␣␣␣␣␣␣␣␣-Immediatecontext:Fulldetailsofadjacentagents
␣␣␣␣␣␣␣␣-Nearbycontext:Keydecisionsfromagents2-3stepsaway
␣␣␣␣␣␣␣␣-Distantcontext:Briefsummariesofagents4+stepsaway
␣␣␣␣␣␣␣␣Considerhowinformationanderrorsmighthavepropagatedacrosstheentireconversationwhenmakingyourargument.

### v2

@@ Abstract
Error attribution in Large Language Model (LLM) multi-agent systems presents a significant challenge in debugging and improving collaborative AI systems. Current approaches to pinpointing agent and step level failures in multi-agent interaction traces—whether using all-at-once evaluation, step-by-step analysis, or binary search—fall short when analyzing complex patterns, struggling with both accuracy and consistency. We present ECHO (Error attribution through Contextual Hierarchy and Objective consensus analysis), a novel algorithm that combines hierarchical context representation, objective analysis-based evaluation, and consensus voting to improve error attribution accuracy. Our approach leverages a positional-based leveling of contextual understanding while maintaining objective evaluation criteria, ultimately reaching conclusions through a consensus mechanism. Experimental results demonstrate that ECHO outperforms existing methods across various multi-agent interaction scenarios, showing particular strength in cases involving subtle reasoning errors and complex interdependencies. Our findings suggest that leveraging these concepts of structured, hierarchical context representation combined with consensus-based objective decision-making, provides a more robust framework for error attribution in multi-agent systems.
**footnotetext: Equal Contribution.
@@ 1 Introduction
Multi-Agent Systems (MASs) have demonstrated remarkable performance across use-cases such as coding [6], medical QA [7], and financial decision-making [8]. However, their multi-step nature makes them vulnerable to compounding errors, where early mistakes amplify through subsequent steps and can derail the entire system. As a result, identifying the initial error’s source—both agent and step—becomes crucial for mitigating such failures and improving these systems.
@@ 2.1 Evaluation of LLM Agents
Much research has focused on automated LLM evaluation across various tasks [13, 14, 15], with recent efforts shifting toward assessing the agentic abilities of LLMs. AgentBench [16] tests task completion across environments like online shopping and householding, while MLAgentBench [17] evaluates LLM agents’ abilities to perform machine learning experimentation. Similarly, AssistantBench [18] features realistic, time-intensive web tasks requiring multi-turn interactions, while Agent-as-a-Judge [19] leverages LLM agents to evaluate other agent peers through tool interactions. Finally, ToolFormer [20], AnyTool [21], and ToolBench [22] benchmark LLMs’ API-calling capabilities via tool-use. As LLM agent interactions evolve, evaluation frameworks now address multi-agent scenarios: MultiAgentBench [23] measures coordination and competitive dynamics, while SwarmBench [24] assesses swarm intelligence through coordination tasks like foraging and flocking.
@@ 2.2 Error Attribution
While LLM evaluation typically measures task completion and output quality, error attribution focuses on systematically classifying error types and patterns to understand fundamental failure modes [25, 26, 27, 28]. ReaLMistake [29] assesses LLMs’ error attribution capabilities in generated output, while SynCheck [30] leverages decoding dynamics to verify sentence trustworthiness and backtracks the unfaithfulness of generated sentences. Self-Backtracking [31] trains LLMs to localize and correct their own reasoning errors through supervised fine-tuning. Furthermore, Process Reward Models (PRMs) [32] evaluate intermediate reasoning step correctness through process annotation, and ProcessBench [33] measures PRMs’ step-wise evaluation capabilities. These approaches focus solely on single-agent error analysis, whereas the Who&When dataset [9] and ECHO extends error attribution to multi-agent settings by identifying both the responsible agent and erroneous step.
@@ 3.1 Hierarchical Context Representation
Hierarchical representation operates at 4 levels $L_{1}$ through $L_{4}$, as seen in Appendix A.1 and A.3, to extract key information from agent / step interactions (via regex pattern matching). While LLMs could potentially be used for more flexible content extraction, the computational overhead would be significant—a trade-off we explore further in Section 4.3. Its implementation employs specialized content extraction mechanisms for each layer $C_{i}$, as seen below:
@@ 3.2 Objective Analysis
This approach prevents echo-chamber effects in attribution through intentional perspective diversity. While this was implemented in ECHO through role-specific prompts and temperatures, the framework accommodates other diversification methods (different models, LLM configurations, or reasoning strategies)—all serving the core principle of bias mitigation through analytical diversity.
@@ 3.3 Consensus Voting
The consensus mechanism processes 3 key components from each analysis $A_{t}^{j}$ (where $t\in\{\text{agent},\text{step}\}$): (1) Primary Conclusions—that considers core attribution decisions including the attribution type, responsible agent(s), and specific error step; (2) Agent Evaluations—that provides assessments of each agent’s error likelihood and supporting evidence; and (3) Alternative Hypotheses—that considers secondary attribution possibilities that maintain reasonable likelihood.
@@ 4.2.1 Performance of ECHO
Step-Level Attribution: Although ECHO still outperforms prior methods with statistical significance, step-level precision still proves challenging, with exact matches at 27–28%. However, accuracy improves significantly when accounting for step tolerance. With ±3 steps, accuracy reaches 42.1% for hand-crafted dataset with ground truth; and when further extended to ±5 steps, accuracy reaches 61.4%—again, maintaining superior performance over baselines.
@@ 4.3 Impact of ECHO Components Via Ablation
Hierarchical context shows clear advantages over fixed context windows within the context-aware framework. Shifting from fixed ±1 step (I1) to hierarchical context (I2) yields significant improvements in hand-crafted datasets: +16.1% in agent-level accuracy (P < 0.05) and +1.9% in step-level accuracy (not significant), as shown in Table 2. A similar trend is exhibited for the algorithm-generated dataset, with a +6.2% in agent-level accuracy and +3.8% in step-level accuracy. This validates graduated context preservation, where detail decreases with distance from target steps. While analysis for the hand-crafted dataset was limited to shorter traces due to computational constraints with context-aware agents, the consistent accuracy improvements highlight hierarchical context’s value—a benefit fully realized when combined with objective analysis.
@@ A.2 Fixed-Window Context
Extract agent contexts from conversation history.
Each context includes the previous agent, current agent, and next agent.
Args:
conversation_history: List of conversation turns with agent information
Returns:
List of tuples containing (prev_agent, current_agent, next_agent) for each agent
@@ A.3 Hierarchical Context Extraction
Extract key decision or main point from agent content using regex patterns.
Args:
agent_content: The full content of the agent
max_words: Maximum words in the extracted key decision
context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
Returns:
Key decision or main point from the agent’s content
@@
Create a brief summary of agent content using regex patterns.
Args:
agent_content: The full content of the agent
max_words: Maximum words in the summary
context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
Returns:
Brief summary of the agent’s content
@@
Extract milestone-based information from agent content using regex patterns.
This provides a higher level of abstraction than brief summaries for distant contexts.
Args:
agent_content: The full content of the agent
max_words: Maximum words in the extracted milestones
context_type: Type of context to focus on (handoff, decision_quality, error_propagation, general)
Returns:
Milestone-based information from the agent’s content
@@
Extract hierarchical agent contexts from conversation history.
Uses graduated detail levels based on distance from current agent.
Args:
conversation_history: List of conversation turns with agent information
dataset_name: Name of the dataset being processed (affects how agent info is extracted)
Returns:
List of dictionaries containing hierarchical context for each agent
@@ A.4 Context Step Aware Agent
Context-Aware Step Agent that analyzes an agent in the context of its previous and next agents
to argue why the error happened in this agent’s step.
@@
Initialize the Context-Aware Step Agent.
Args:
model_id: The model ID to use for the agent
temperature: The temperature to use for generation
@@
You are a Context-Aware Step Agent that analyzes an agent’s actions in the context of the previous and next agents.
Your task is to argue why the error happened in YOUR agent’s step.
Your task:
1. Analyze what information was received by your agent from the previous agent (if any)
2. Analyze what information was generated by your agent
3. Analyze how your agent’s output affected the next agent (if any)
4. Make a strong argument for why YOUR AGENT caused the final wrong answer, using the ground truth as evidence
Input:
- Ground Truth: [GROUND_TRUTH]
- Final Answer: [FINAL_ANSWER]
- Agent Context: Information about the previous, current, and next agents
Output your response with the following clear section headers:
## Purpose:
Describe the purpose of this agent step - what was this agent trying to accomplish?
## Assumptions and Information:
List the assumptions and information this agent was given from the previous agent or context.
## Errors:
Describe what this agent did wrong (if anything). Be specific about any mistakes, misunderstandings, or incorrect reasoning.
## Evidence:
Provide evidence from the ground truth that supports your error attribution. Explain how this agent’s actions directly led to the wrong final answer.
Remember: You must argue that YOUR agent caused the error. Be persuasive and use evidence.
@@ 3. Analyze how your agent’s output affected the next agent (if any)
Analyze an agent in the context of its previous and next agents and generate an argument
for why this agent caused the error.
Args:
step_id: The ID of the step (e.g., "step_1")
prev_agent: The previous agent (or None if first agent)
current_agent: The current agent being analyzed
next_agent: The next agent (or None if last agent)
ground_truth: The ground truth answer
final_answer: The final answer given
query: The original query/question
Returns:
JSON argument for why this agent caused the error
@@
Original Query: {query}
{ground_truth_section}
Final Answer: {final_answer}
Agent Context:
{agent_context}
Please analyze this agent in the context of the previous and next agents, and provide a strong argument for why THIS agent caused the final wrong answer.
Use the section headers specified in your instructions (Purpose, Assumptions and Information, Errors, Evidence).
@@
Analyze an agent using hierarchical context and generate an argument
for why this agent caused the error.
Args:
step_id: The ID of the step (e.g., "step_1")
hierarchical_context: Dictionary containing hierarchical context information
ground_truth: The ground truth answer
final_answer: The final answer given
query: The original query/question
Returns:
JSON argument for why this agent caused the error
@@
Original Query: {query}
{ground_truth_section}
Final Answer: {final_answer}
Agent Context:
{agent_context}
Please analyze this agent in the hierarchical context of the entire conversation, and provide a strong argument for why THIS agent caused the final wrong answer.
Use the section headers specified in your instructions (Purpose, Assumptions and Information, Errors, Evidence).
Note: You now have access to the full conversation context at different detail levels:
- Immediate context: Full details of adjacent agents
- Nearby context: Key decisions from agents 2-3 steps away
- Distant context: Brief summaries of agents 4+ steps away
Consider how information and errors might have propagated across the entire conversation when making your argument.